OpenStudy (anonymous):

Find the circumference of a circle with area 25p m2.

OpenStudy (anonymous):

p m2 mean..?

OpenStudy (anonymous):

the formula for circumfrence is =2*pi*r here r=radius

OpenStudy (anonymous):

oh...so compare it with the area of circle i.e. pi*r^2 to find r..as pi*r^2=25*pi m2 =>r^2=25 m2 =>r=5m now circumfrnec is =2*pi*r =< C=2*pi*5m=10pi=10*3.14=31.4m